# Univariate-Guided Sparse Regression

### v8

@@ 8.2 Real data
We compared least squares and uniReg on the regression datasets from the UCI database. There were 23 datasets in all. Some had multiple targets: we treated each target separately and averaged the results. We randomly sampled each dataset into a $70/30$ training/test split, and report the test set MSE and support in Figure 10. One dataset had s small sample ($n=60$) which produced a very large MSE from least squares; we omit this from the plot for visibility.
@@
The two methods perform similarly in MSE; but uniLasso has much smaller support, averaging about $1/3$ that of the least squares support (which has the number of features $p$).
@@ 9 Real data examples for uniLasso
10 More examples
10.1 UniLasso and adversarial attacks
In machine learning terminology, an adversarial attack creates an altered input that leads to a large error in the model’s prediction. In image recognition problems, the attack also produces an image that is imperceptible to humans, In statistical terms, an adversarial attack creates features in the test data that come from a different distribution than the training data,
A common way of creating an adversarial attack is to add noise in the direction of the gradient of the objective function (e.g. ?). We investigate this here in the context of the proteomics example of Table 1. We carried out the following process:
(a) Randomly split the data into a 70/30 train test split.
(b) Fit lasso and uniLasso to the training data
(c) Add varying levels of Gaussian noise to the test features in the direction of the (sub) gradient of each model from (b), and compute the test set AUC.
We repeated this 200 times and plot the results in Figure 12.
Figure 12: Results of adversarial attacks in the proteomics example of Table 1: test set AUC
We see that uniLasso is less affected by the attack than the lasso, presumably because the univariate guidance disallows sign changes and the individual features that are selected are stronger on their own.
10.2 Multiclass setting and small round blue cell cancers
The multinomial approach to the multiclass problem does not generalize easily to UniLasso. The reason is that the coefficients for each class are only determined up to a shift, and hence a non-negativity constraint does not make sense. Instead we take a “one-versus rest” (OVR) approach, applying UniLasso binary classifier, to classify each class from the others. Then we take the class with the highest probability as the predicted class.
We applied this to a dataset on childhood cancer, with 63 training and 25 test samples in 4 classes. (?) With the original train/test split we get zero test errors with glmnet. We repeated the train/test split 50 times and obtained these results:
mean errors out of 25 sd mean(supp) sd(supp)
@@ 10.2 Multiclass setting and small round blue cell cancers
10.3 A GWAS example
The data here come from the Kaggle site
It is described as a simulated dataset comparable to the Illumina 650K human array, for SNP genotyping. There are 1000 individuals, and Genotypes (0, 1 or 2) for 482906 markers For illustration, we selected the 10,000 markers with highest variance and divided the data into training and test sets each of size 500. The target measure is quantitative.
Figure 13: GWAS example: CV and test set R-squared curves
In Figure 13 we see that UniLasso delivers a model with slightly higher test $R^{2}$ using fewer SNPs. Table 8 summarizes the signs of the Lasso and UniLasso coefficients in final chosen models. Interestingly, ignoring the zeroes, there are no sign disagreements!
uni$<$0 uni=0 uni$>$0
lasso$<$0 33 13 0
lasso=0 1 9901 0
lasso$>$0 0 13 39
Table 8: GWAS example: signs of Lasso and UniLasso coefficients in final chosen models.
@@ 11 The use of external univariate scores
Consider the setting where we have our training set T and also external data E from the same domain (e.g. disease), We assume that E does not contain raw data but only summary results. Specifically, E contains just univariate coefficients and standard errors for each feature. This setting occurs fairly often in bio-medical settings where investigators are not willing to share their raw data, but do publish and share summary results.
@@
We demonstrate here how uniLasso can make productive use of external scores. The idea is to use the univariate coefficients from E, rather than computing LOO estimated from the training set. Then in step 2 of uniLasso, we proceed as usual.
@@
To investigate this scheme, we generated data from our ‘homecourt” setting of Section 7 ($n=300,p=1000,SNR=1.5$), as well as 1000 extra samples from the same distribution. Figure 14 shows the test set MSE over 200 simulations from the following strategies:
  • the test set MSE for the lasso, and uniLasso, both applied just to the training set
@@
  • uniLasso+$m$, where the univariate scores derived from the $m$ extra samples are used in place of the LOO estimates
  • uniLasso+all- uniLasso trained on the combined 1300 samples.
The last estimate is there for reference: it is not available without access to the raw extra data.
We see that adding the scores from just 100 external observations improves the performance of uniLasso. Perhaps surprisingly, we don’t see any improvement with additional scores. The rightmost boxplot shows that access to the full external data (not just the scores) would lead to further improvement.
Figure 14: Results from an experiment examining the use of external data with uniLasso. See text for details.
@@ 12 Does CV work in our two stage UniLasso procedure?
In the UniLasso algorithm we use the target $y$ in both steps, but for (significant) computational convenience, we only apply cross-validation in the 2nd step. Thus one should be concerned that cross-validation may not perform well here. In the simulations of Section 7 we used this form of cross-validation to choose the $\lambda$ tuning parameter and it seems to have done a reasonably good job.
A related question is whether the error reported by CV is a good estimate of test error, especially at the selected value of $\lambda$. Figure 15 shows the CV error at the selected value of $\lambda$ versus the test error, for the Lasso (left) and UniLasso (right). We see that both CV estimates are quite good, with both over-estimating the test error in a few cases. It seems clear that our use of cross-validation in UniLasso is not a major concern..
Figure 15: CV error at the selected value of $\lambda$ versus the test error, for the Lasso (left) and UniLasso (right). Each datapoint represents one of our simulated or real datasets studied earlier. Both CV estimates are quite good, with both over-estimating the test error in a few cases.
@@ 13 UniLasso for GLMs and the Cox Survival Model
    Fit the $p$ univariate GLMS, and produce the linear predictor functions $\hat{\eta}_{j}(x_{j})$. Compute the LOO predictions $\hat{\eta}_{j}^{-i}$ for the $n$ training observations for each of these.
@@ 14.1 Least Squares
We first discuss the computations for squared-error loss. For feature $j$ we have to fit a univariate linear model, and then compute its LOO predictions. We are required to fit the model $\eta_{j}(x_{j})=\beta_{0j}+\beta_{j}x_{j}$ using the data $\{x_{ij},y_{i}\}_{1}^{n}$. This task is simplified if we standardized $x_{j}$ to have mean zero and unit variance: $z_{j}=(x_{j}-\bar{x_{j}})/s_{j}$, where $\bar{x}_{j}=\frac{1}{n}\sum_{i=1}^{n}x_{ij}$, and $s_{j}^{2}=\frac{1}{n}\sum_{i=1}^{n}(x_{ij}-\bar{x}_{j})^{2}$. Then the least squares estimates using $z_{j}$ are $(\hat{\delta}_{0j},\;\hat{\delta}_{j})=(\bar{y},\;\frac{1}{n}\sum_{i=1}^{n}z_{ij}y_{i})$. These are mapped back to least squares estimates for $x_{j}$ via $(\hat{\beta}_{0j},\;\hat{\beta}_{j})=(\hat{\delta}_{oj}-\hat{\delta_{j}}\bar{x}_{j}/s_{j},\;\hat{\delta}_{j}/s_{j})$.
This standardized form is particularly useful for computing the LOO fits, since fits are invariant under affine transformations of the features. We have the classical formula for the LOO residual in linear regression
@@
where $H_{ii}$ is the $i$th diagonal entry of the hat matrix. It is straightforward to show that using the $z_{j}$ gives $H_{ii}=1/n+z_{ij}^{2}/n$. Now simple algebra backs out an expression for $\hat{\eta}_{j}^{-i}$.
@@ 14.2 Generalized Linear Models
Once again we can fit all the univariate GLMs using matrix operations at the same time — i.e. we perform a each IRLS step simultaneously for all $p$ univariate GLMS. The expressions are only slightly more complex than in the unweighted case.
What about the LOO fits $\hat{\eta}_{j}^{-i}$ for each univariate GLM? Here we make an approximation, as recommended by (?), which amounts to using the final weighted least squares IRLS iteration when fitting the models $\hat{\eta}_{j}(x_{j})$. A simple formula is available there, similar to what we got before, except a bit more detailed to accommodate observation weights. Again we can use Hadamard matrix operations to do this simultaneously for all $i$ and $j$.
For the Cox model, we make a further approximation. The implied weight matrix in IRLS is not diagonal. We simply use the diagonal which leads to an approximate Newton algorothim. Note that for the Cox model the intercept is always zero.
@@ 14.3 Software in R
We have written an R function cv.UniLasso, whose arguments mirror cv.glmnet. This function does the following, currently for "binomial", "gaussian" and "cox" families. It does all the work outlined above, and performs the second stage cv.glmnet using the (approximate) LOO $\hat{\eta}_{j}^{-i}$ as features. A cv.glmnet object contains information about suggested values of $\lambda$, as well as a fitted glmnet model fit to all the data with solutions at all values of $\lambda$. This object is referenced when one predicts from a cv.glmnet object. In this case we make sure that the coefficients that are stored on this objects for each $\lambda$ are the collapsed versions as in (10).
@@
    lower.limits = 0: this default choice guarantees theat the $\hat{\theta}_{j}$ are all non-negative. We recommend this in order to get a sparser and more interpretable model.

### v9

@@ 9.2 Real data
We compared least squares and uniReg on the regression datasets from the UCI database [^4] https://archive.ics.uci.edu/datasets. There were 23 datasets in all. Some had multiple targets: we treated each target separately and averaged the results. We randomly sampled each dataset into a $70/30$ training/test split, and report the test set MSE and support in Figure 8. One dataset had a small sample ($n=60$) which produced a very large MSE from least squares; we omit this from the plot for visibility.
@@
The two methods perform similarly in MSE; but uniReg has much smaller support, averaging about $1/3$ that of the least squares support (which is the number of features $p$).
9.3 Theory for uniReg without LOO
Our data consists of $(X_{i},Y_{i})$, $i=1,\ldots,n$, where
$Y_{i}=\beta_{0}+X_{i}^{T}\beta+\epsilon_{i},$
where $\beta_{0}\in\mathbb{R}$, $\beta=(\beta_{1},\ldots,\beta_{p})\in\mathbb{R}^{p}$, $X_{i}$ are i.i.d. $N_{p}(0,\Sigma)$ random vectors, where $\Sigma$ is a positive definite matrix with minimum eigenvalue $\lambda_{0}$ and maximum eigenvalue $\lambda_{1}$, and $\epsilon_{i}$ are i.i.d. $N(0,\sigma^{2})$ random variables for some $\sigma>0$. Let $r:=p/n$. We assume that $r\in(0,r_{0})$ for some $r_{0}<1$.
Let $\hat{\beta}_{0}^{\mathrm{OLS}}$ and $\hat{\beta}^{\mathrm{OLS}}$ denote the OLS estimates of $\beta_{0}$ and $\beta$. Let $\hat{\beta}_{0}^{\mathrm{UR}}$ and $\hat{\beta}^{\mathrm{UR}}$ denote the uniReg estimates of $\beta_{0}$ and $\beta$ without leave-one-out, defined as follows. First, compute the univariate coefficients $\hat{\beta}^{\mathrm{uni}}_{j}$, $j=1,\ldots,p$, as the coefficient of $X_{j}$ when $Y$ is regressed solely on $X_{j}$ with an intercept term. The estimates $\hat{\beta}_{0}^{\mathrm{UR}}$ and $\hat{\beta}^{\mathrm{UR}}$ are obtained by minimizing
$\sum_{i=1}^{n}(Y_{i}-a-X_{i}^{T}b)^{2}$
over $a\in\mathbb{R}$ and $b\in\mathbb{R}^{p}$ subject to the constraint that $b_{j}\hat{\beta}^{\mathrm{uni}}_{j}\geq 0$ for $j=1,\ldots,p$. Let $\beta^{\mathrm{uni}}_{0,j}$ and $\beta^{\mathrm{uni}}_{j}$ be the population univariate coefficients; that is,
$\rm E(Y_{i}|X_{i,j})=\beta^{\mathrm{uni}}_{0,j}+\beta^{\mathrm{uni}}_{j}X_{i,j}.$
Define three vectors $\mu,\hat{\mu}^{\mathrm{OLS}},\hat{\mu}^{\mathrm{UR}}\in\mathbb{R}^{n}$ as
$\mu_{i}:=\rm E(Y_{i}|X_{i})=\beta_{0}+X_{i}^{T}\beta,$
$\hat{\mu}^{\mathrm{OLS}}_{i}:=\hat{\beta}_{0}^{\mathrm{OLS}}+X_{i}^{T}\hat{\beta}^{\mathrm{OLS}},$
$\hat{\mu}^{\mathrm{UR}}_{i}:=\hat{\beta}_{0}^{\mathrm{UR}}+X_{i}^{T}\hat{\beta}^{\mathrm{UR}},$
for $i=1,\ldots,n$. Let $q$ be the number of $j$ such that $\beta_{j}=0$ and $\beta^{\mathrm{uni}}_{j}\neq 0$. The following theorem shows that if $\beta_{j}$ and $\beta^{\mathrm{uni}}_{j}$ have the same sign for all $j$, then $\rm E\|\mu-\hat{\mu}^{\mathrm{UR}}\|^{2}$ is less than or equal to $\rm E\|\mu-\hat{\mu}^{\mathrm{OLS}}\|^{2}$ minus a constant times $q$.
Theorem 4.
Suppose that $\beta^{\mathrm{uni}}_{j}\beta_{j}\geq 0$ for each $1\leq j\leq p$. Let $\delta$ be the minimum of $|\beta^{\mathrm{uni}}_{j}|$ over $1\leq j\leq p$ such that $\beta^{\mathrm{uni}}_{j}\neq 0$. Then there are positive constants $C_{0},C_{1}$ depending only on $\lambda_{0}$, $\lambda_{1}$, $\delta$, $r_{0}$, and $\sigma$ such that if $n\geq C_{0}$, then
$\rm E\|\mu-\hat{\mu}^{\mathrm{UR}}\|^{2}\leq\rm E\|\mu-\hat{\mu}^{\mathrm{OLS}}\|^{2}-C_{1}q.$
The theorem is proved in Appendix C. We have not been able to prove the LOO version of uniReg (that is, the actual version that we have proposed). We leave it as an open question.
@@ 10 Real data examples for uniLasso
11 Multiclass models
Here we consider a classification problem with more than two classes. The multinomial model is natural for the multiclass problem, but does not generalize easily to uniLasso. The reason is that for each feature there are coefficients for each class, so it is not clear how to replace the feature with a LOO version. Furthermore the coefficients per class are only determined up to a shift, and hence a non-negativity constraint does not make sense. Instead we take a “one-versus rest” (OVR) approach, where the uniLasso binary classifier is used to classify each class from the others. Hence each class gets a predicted probability, and with OVR one classifies to the class with the highest probability as the predicted class.
We applied this to a dataset on childhood cancer, with 63 training and 25 test samples in 4 classes (Khan et al., 2001). With the original train/test split we get zero test errors with glmnet. We repeated the train/test split 50 times and obtained these results:
Ave # Errors SD Ave Support SD
@@ 11 Multiclass models
Table 7: Results for the four-class cancer example. Overall uniLasso OVR makes the fewest misclassification errors averaged over the 50 train-test splits. Of the 50 train(63)/test(25) splits, 41 made zero errors on the 25 test data points, and 9 made 1 error, averaging 0.18. Lasso OVR is sligthly worse, and lasso multinomial makes almost one error on average per train/test split. On the other hand lasso multinomial yields the sparsest model, since the other two select features separately for each class.
@@ 12 The use of external univariate scores
Consider the setting where we have our training set T and also external data E from the same domain (e.g. disease), We assume that E does not contain raw data but only summary results. Specifically, E contains just univariate coefficients and standard errors for each feature. This setting occurs fairly often in biomedical settings where investigators are not willing to share their raw data, but do publish and share summary results.
@@
We demonstrate here how uniLasso can make productive use of external scores. The idea is to use the univariate coefficients from E, rather than computing the LOO estimates from the training set. Then in step 2 of uniLasso, we proceed as usual.
@@
To investigate this scheme, we generated data T with $n=300$, $p=1000$, $\mbox{SNR}=1.5$, with feature covariance AR(1), with $\rho=0.8$ and the non-zero regression coefficients distributed as $U(0.5,2)$, and positioned on every other feature $(1,3,\ldots 99)$. We also generated 600 extra samples from the same distribution to serve as an external data set E, and in addition a large test set. Figure 10 shows the test-set MSE over 200 simulations from the following strategies:
  1. 1.
    lasso and uniLasso, both applied just to the training set;
  2. 2.
@@
    uniLasso+$m$ (green) where the univariate scores are derived from $m$ extra samples from E, and are used in place of the LOO estimates;
  3. 3.
    unilLasso (red) on all $300+m$ samples;
Figure 10: Results from an experiment examining the use of external data with uniLasso. The initial training set has 300 observations and we generated 600 additional samples. Shown in blue are the test set MSE for the lasso, and uniLasso, both applied just to the training set; uniLasso+$m$ (green), where the univariate scores derived from the $m$ extra samples are used in place of the LOO estimates, and unilLasso (red) on $300+m$ samples for $m=50,100\ldots 300$.
We see that adding the scores from just 50 external observations does slightly worse than uniLasso on the original data, but builds sparser models than lasso at no cost in test MSE. If instead, we use the extra 50 samples as in strategy 3, the performance improves.
If we increase $m$ to 100, both strategies 2 and 3 improve, but uniLasso using the augmented data performs better each time.
So it appears strategy 2 eventually (as we increase $m$) outperforms lasso and uniLasso both fit on only the original 300 observations. But if we have the raw external data, strategy 3 is the clear winner.
This is one simulation scenario, and it is of course possible that in others the pattern may not be as clear.
13 A uniLasso polish
As we have seen earlier, the uniLasso procedure can sometimes perform considerably worse than the lasso, especially in cases where its sign constraint is too restrictive. An example is the “counter-example” problem in Section 8 , where two positively correlated features have different coefficient signs in the population multivariate model. While problematic, the CV estimate of error will alert the user that uniLasso is not working well in a given problem. In this section, we propose a simple “post-processing” of the uniLasso solution that can help to diagnose shortcomings of uniLasso and remedy them.
The idea is very simple: we run the usual lasso (with-cross-validation) to the residual from the uniLasso fit, and then “stitch” together the two solutions. Here is the algorithm in detail, using the glmnet R package terminology:
  1. 1.
    Run cv.uniLasso on $X,y$ , and extract the fitted linear predictor $\hat{y}$ at $\hat{\lambda}_{min}.$
  2. 2.
    Apply cv.glmnet on $X,y$ , with offset $\hat{y}$, giving final predictions $\hat{y}_{g}$.
  3. 3.
    “Stitch” together the two solutions to produce a single path, starting at the original uniLasso solution.
Note that in the Gaussian model the use of the offset is equivalent to applying the lasso to the residual in step 2; expressed as an offset it allows application of the idea to other GLM families.
Figure 11: UniLasso polish applied to the counter-example problem of Section 8. The left panel shows the uniLasso solution path, with the dashed vertical line indicating the solution chosen by cross-validation. The right panel shows the polish.uniLasso solution path, which starts at the uniLasso, and the vertical dashed line indicates the model chosen by cross-validation. The red curves are the support features $x_{1}$ and $x_{2}$.
Figure 11 shows the results of the uniLasso polish applied to the counter-example problem of Section 8. There are two support features ($x_{1}$ and $x_{2}$). In the left panel uniLasso enters just feature $x_{1},$ because $x_{2}$ has a positive univariate coefficient but a negative multivariate coefficient. The right panel shows the “polished” path. Via the offset it starts at the uniLasso solution (chosen by CV), and immediately enters $x_{2}$, and eventually growing the coefficients of both $x_{1}$ and $x_{2}$. In the process, the model lets in some of the noise variables.
In Section 8 we included the uniLasso polish in all of the settings, and its performance overall is excellent.
@@ 14 Does CV work in our two stage uniLasso procedure?
In the uniLasso algorithm we use the target $y$ in both steps, but for (significant) computational convenience, we only apply cross-validation in the 2nd step. Thus one should be concerned that cross-validation may not perform well here. In the simulations of Section 8 we used this form of cross-validation to choose the $\lambda$ tuning parameter, and it seems to have done a reasonably good job.
Figure 12: CV error at the selected value of $\lambda$ versus the test error, for the lasso (left) and uniLasso (right). Each datapoint represents one of our simulated datasets with $p>N$ studied earlier. Both CV estimates track the test error reasonably well.
A related question is whether the error reported by CV is a good estimate of test error, especially at the selected value of $\lambda$. Figure 12 shows the CV error at the selected value of $\lambda$ versus the test error, for the lasso (left) and uniLasso (right). We see that both CV estimates are quite good. It seems clear that our use of cross-validation in uniLasso is not a major concern.
Finally, a referee asked whether a reason for uniLasso’s strong performance was its ability to estimate the the CV tuning parameter $\lambda$. For the car price data, Figure 13 shows the ratio of the achieved test error using CV divided by the minimum test error over the $\lambda$ path, for each of the 50 train/test splits. We see that both lasso and uniLasso do an (equally) good job of estimating the optimal $\lambda$. The average ratio $\lambda_{cv}/\lambda_{opt}$ was 0.87 and 0.38 for the lasso and uniLasso, respectively.
Figure 13: Car price data: ratio of the achieved test error using CV divided by the minimum test error over the $\lambda$ path.
@@ 15 UniLasso for GLMs and the Cox survival model
    Fit the $p$ univariate GLMS, and produce the linear predictor functions $\hat{\eta}_{j}^{i}=\hat{\beta}_{0j}+\hat{\beta}_{j}x_{ij}$. Compute the LOO predictions $\hat{\eta}_{j}^{-i}$ for the $n$ training observations for each of these.
@@ 16.1 Least squares
We first discuss the computations for squared-error loss. For feature $j$ we have to fit a univariate linear model, and then compute its LOO predictions. We are required to fit the model $\eta_{j}(x_{j})=\beta_{0j}+\beta_{j}x_{j}$ using the data $\{x_{ij},y_{i}\}_{1}^{n}$. This task is simplified if we standardized $x_{ij}$ to have mean zero and unit variance: $z_{ij}=(x_{ij}-\bar{x_{j}})/s_{j}$, where $\bar{x}_{j}=\frac{1}{n}\sum_{i=1}^{n}x_{ij}$, and $s_{j}^{2}=\frac{1}{n}\sum_{i=1}^{n}(x_{ij}-\bar{x}_{j})^{2}$. Then the least squares estimates using the $z_{ij}$ are $(\hat{\delta}_{0j},\;\hat{\delta}_{j})=(\bar{y},\;\frac{1}{n}\sum_{i=1}^{n}z_{ij}y_{i})$. These are mapped back to least squares estimates for $x_{ij}$ via $(\hat{\beta}_{0j},\;\hat{\beta}_{j})=(\hat{\delta}_{oj}-\hat{\delta_{j}}\bar{x}_{j}/s_{j},\;\hat{\delta}_{j}/s_{j})$.
This standardized form is particularly useful for computing the LOO fits, since the fits are invariant under affine transformations of the features (so we can use the $z_{ij}$ instead of the $x_{ij}$). We have the classical formula for the LOO residual in linear regression
@@
where $H_{ii}$ is the $i$th diagonal entry of the hat matrix. It is straightforward to show that using the $z_{ij}$ gives $H_{ii}=1/n+z_{ij}^{2}/n$. Now simple algebra backs out an expression for $\hat{\eta}_{j}^{-i}$.
@@ 16.2 Generalized linear models
Once again we can fit all the univariate GLMs using matrix operations at the same time — i.e. we perform each IRLS step simultaneously for all $p$ univariate GLMS. The expressions are only slightly more complex than in the unweighted case.
What about the LOO fits $\hat{\eta}_{j}^{-i}$ for each univariate GLM? Here we make an approximation, as recommended by Rad & Maleki (2020), which amounts to using the final weighted least squares IRLS iteration when fitting the models $\hat{\eta}_{j}(x_{j})$. A simple formula is available there, similar to what we got before, except a bit more detailed to accommodate observation weights. Again we can use Hadamard matrix operations to do this simultaneously for all $i$ and $j$.
For the Cox model, we make a further approximation, since the implied weight matrix in IRLS is not diagonal. We simply use the diagonal which leads to an approximate Newton algorithm. Note that for the Cox model the intercept is always zero.
@@ 16.3 Software in R
We provide a R package uniLasso for fitting the models described in this paper. It mirrors the behavior of the glmnet package for fitting lasso models. The function cv.uniLasso has arguments that include all those for cv.glmnet plus a few extras. This function does the following, currently for "binomial", "gaussian" and "cox" families. It does all the work outlined in Sections 16.116.2, and performs the second stage cv.glmnet using the (approximate) LOO $\hat{\eta}_{j}^{-i}$ as features. A cv.glmnet object contains information about suggested values of $\lambda$, as well as a fitted glmnet model fit to all the data with solutions at all values of $\lambda$. This object is referenced when one predicts from a cv.glmnet object. In this case we make sure that the coefficients that are stored on this objects for each $\lambda$ are the collapsed versions as in (9).
@@
    lower.limits = 0: this default choice guarantees that the $\hat{\theta}_{j}$ are all non-negative. We recommend this in order to get a sparser and more interpretable model.
@@
In addition we have uniReg and cv.uniReg for fitting the uniReg model, as well as ci.uniReg for computing bootstrap intervals for each coefficient. The function polish.uniLasso does the polishing as described in Section!13. We also provide simulation functions for generating the data as used in this paper.